OpenStudy (anonymous):

Subtract. simplify by collecting like radical term if possible 8sqrt20 -6sqrt5 I got 10sqrt5 is that right????????????

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[10\sqrt{5}\]

OpenStudy (mathteacher1729):

\[8\sqrt{20} - 6\sqrt{5}\] Let's look at the square root of 20 first. 20 = 4*5 and since 4 is a perfect square we can write: \[\sqrt{20} = \sqrt{4\cdot 5}=\sqrt{4}\cdot\sqrt{5}=2\sqrt{5}\] So instead of writing \[\sqrt{20}\] write \[2\sqrt{5}\] and now we have: \[8\sqrt{20}=8\cdot (2\sqrt{5})=16\sqrt{5}\]. And \[16\sqrt{5}-6\sqrt{5} =10\sqrt{5}\] :)

OpenStudy (mathteacher1729):

Argh, I must learn to type faster. :-p

OpenStudy (anonymous):

No your ok, thank you for your help